OpenStudy (anonymous):

Is this the simplified version of √20? 2√5

OpenStudy (agent0smith):

Yes it is. \[\sqrt{20} = \sqrt{5 \times 4} = 2\sqrt5\]

OpenStudy (anonymous):

Thanks just checking :)